solve the quadratic equation 2xpower 2 +ax-a power 2=0 for x.

Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
2x^{2}+ax-a^{2}=0
2x^{2}+2ax-ax-a^{2}=0
2x(x+a)-a(x+a)=0
(2x-a)(x+a)=0
x = \frac{a}{2}, -a
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty